# Frostman random variables, entropy inequalities, and applications

### v3

@@ 7 Examples
In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Here measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
$P(X\in E)=\frac{\mathcal{L}^{1}(E\cap A)}{\mathcal{L}^{1}(A)}$
for any Borel set $E$. In the following, we will proof several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
@@
Let us consider the probability that $X$ takes a value in any ball $B(x,r)$. Calculation reveals that
@@
One can verify that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$. Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$. ∎
@@
$H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1),$
and
$H(X^{\prime}+Y^{\prime})\geq H(X^{\prime})=m.$
@@
$H_{n}(X^{\prime 2}+Y^{\prime 2})=\sum\limits_{i=0}^{2^{n+1}}\iota\Big{(}P(D_{n}(X^{\prime 2}+Y^{\prime 2})\quad=i2^{-n})\Big{)}$
$=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big{(}P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big{)}=\sum\limits_{k\in\mathcal{K}}\iota\big{(}2^{-2m}T(k)\big{)},$
@@
On the other hand, one has $T(k)\geq 1$ for all $k\in\mathcal{K}$. So
@@
So, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
@@
and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
@@
Furthermore, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
@@
and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$. So $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman.
@@
One can verify that all the terms on the right-hand side of above formula can be bounded by $2^{m(2s-1)}r^{s}$ for any $1/2\leq s\leq 2$. The random variable $(X,Y)$ is $\left(s,O(2^{m(2s-1))}\right)$-Frostman. ∎
@@
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))$
@@
The rest cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}{1/2}.$
@@ Appendix A Appendix - Existence of Certain Random Variable
So $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e. $(Y^{\prime\prime}|W^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$. Taking $W=(X,Y)$, one gets $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
@@ Acknowledgements
Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working condition.

### v4

@@ 7 Examples
In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have uniform distribution on $A$. In the following, we will prove several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman, whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
@@
Consider the probability that $X$ takes a value in any ball $B(x,r)$. Calculation reveals that
@@
Calculation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$. Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$. ∎
@@
$m=H(X^{\prime})\leq H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1).$
@@
$H_{n}(X^{\prime 2}+Y^{\prime 2})=\sum\limits_{i=0}^{2^{n+1}}\iota\Big(P(D_{n}(X^{\prime 2}+Y^{\prime 2})=i2^{-n})\Big)$
$\qquad=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big(P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big)=\sum\limits_{k\in\mathcal{K}}\iota\big(2^{-2m}T(k)\big),$
@@
On the other hand, one has $T(k)\geq 1$ for all $k\in\mathcal{K}$. Therefore,
@@
Hence, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
@@
and $Y$ has uniform distribution on $0,2^{-2m}]$. Verification gives $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
@@
In addition, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
@@
and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. Calculation reveals that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$. So $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman.
@@
One can check that all the terms on the right-hand side of above formula can be bounded by $2^{m(2s-1)}r^{s}$ for any $1/2\leq s\leq 2$. The random variable $(X,Y)$ is $\left(s,O(2^{m(2s-1))}\right)$-Frostman. ∎
@@
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculations reveal that
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll M(r_{1})M(r_{2})\cdot 2^{4m}$
where
@@
In all cases, we have
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}^{1/2}.$
@@ Appendix A Appendix - Existence of Certain Random Variable
It follows that $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e. $(Y^{\prime\prime}|W^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$. Taking $W=(X,Y)$, one gets $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
@@ Acknowledgements
Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working conditions.